### v2

# CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features

Abstract
Effective deepfake detection tools are becoming increasingly essential to the growing usage of deepfakes in unethical practices. There exists a wide range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism. The 2025 IEEE Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images containing significant class imbalance. The images in the dataset are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection. To this end, we proposed a disjoint set-based multistage training method to address the class imbalance and devised an ensemble-based architecture CAE-Net. Our architecture consists of a convolution- and attention-based ensemble network, and employs three different neural network architectures: EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet transform to capture both local and global features of deepfakes. We visualize the specific regions that these models focus on for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots. Individually, the EfficientNet B0 architecture has achieved 90.79% accuracy, whereas the ConvNeXt and the DeiT architecture have achieved 89.49% and 89.32% accuracy, respectively. With these networks, our weighted ensemble model achieves an excellent accuracy of 94.63% on the validation dataset of the SP Cup 2025 competition. The equal error rate of 4.72% and the Area Under the ROC curve of 97.37% further confirm the stability of our proposed method. Finally, the robustness of our proposed model against adversarial perturbation attacks is tested as well, showing the inherent defensive properties of the ensemble approach.
keywords:
Deepfake detection , ConvNext , Transformer , DeiT , Wavelet , EfficientNet
[…]
[inst1]organization=Department of Electrical and Electronic Engineering,addressline=BUET, city=Dhaka, postcode=1205, country=Bangladesh
[…]
[inst2]organization=School of Data and Sciences,addressline=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
1 Introduction
Deepfakes are media that are edited or created by generative Artificial Intelligence (AI), which do not exist or have not occurred in reality. Although deepfakes have existed for a long time, it took experts, and often entire studios, to generate convincing fake media. Due to the surge of machine learning and generative AI tools, the generation of deepfakes has never been easier and more accessible. Such ease of access has raised serious societal concerns since these are being used to manipulate public opinion, commit fraud, create non-consensual pornography, and attack individuals through defamation or identity threats [1]. Deepfakes are also being used to gain malicious political advantage worldwide [2]. Although deepfake generation tools are widely accessible, detection tools are not nearly as effective and accurate. Thus, effective detection of deepfakes is a task of great importance and has been extensively studied for the past decade.
The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [4] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset. Wolter et al. [5] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [6] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [7]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [8]. However, Abdullah et al. [9] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [10, 11]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [12]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [13]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [13]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [14, 15].
Researchers are not only relying on complex network architectures but also exploring other approaches, such as the Frequency Enhanced Self-Blending Images (FSBI) method for deepfake detection, which utilizes the Discrete Wavelet Transform (DWT) [16]. Zhu et al. introduced a learnable Image Decomposition Module (IDM) that uses a recomposition operation to highlight illumination inconsistencies [17]. It features a multi-level enhancement technique for effective feature recomposition and is trained in the logarithmic domain, with an extensible design [17]. To address challenges with low-quality datasets and inadequate detection performance, Cheng et al. have developed MSIDSnet [18]. This framework utilizes a multi-scale fusion (MSF) module to capture forged facial features and an interactive dual-stream (IDS) module to enhance feature integration across frequency and spatial domains [18]. Another recent detection model by Byeon et al. uses a spatial-frequency joint dual-stream convolutional neural network with learnable frequency-domain filtering kernels [19].
Research gap and our contribution. The existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector. That is, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features and as such, ii) the inability to be up to date with new fake image generation techniques [20]. To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm. Such a diverse dataset poses unique challenges to formulating accurate image detection methods. In our proposed solution for generalized deepfake image detection, we have employed a weighted ensemble approach, CAE-Net, which utilizes three different neural network architectures with different strengths: i) an EfficientNet architecture, ii) a Data Efficient Image transformer model, and iii) a ConvNeXt model. More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result. We note that the EfficientNet-based architecture, due to its depthwise separable convolution and squeeze and excitation block, allows learning fine spatial artifacts and enhances channel-wise attention [21]. Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image. More specifically, the EfficientNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [22]. Finally, the ConvNeXt model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images. We have utilized the “Haar” wavelets due to their notable ability to detect abrupt transitions and edges in the data [23]. A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone. Finally, we test our models against adversarial perturbation attacks, which shows the trade-offs between weighted ensemble and majority voting ensemble methods.
The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, the description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt. Section 3 contains the accuracy results of multiple standard models on the dataset, along with the implementation details and other evaluation metrics of our proposed approach. This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations. Finally, Section 4 draws concluding remarks on the contributions presented in the paper.
2 Proposed Methodology: Weighted Ensemble of Neural Network Models
2.1 Description of the Dataset
As mentioned before, the DFWild-Cup competition has provided a diverse dataset. The dataset contains images from eight publicly available standard datasets designed for the DeepfakeBench evaluation [24]: Celeb-DF-v1 [25], Celeb-DF-v2 [25], FaceForensics++ [26], DeepfakeDetection [27], FaceShifter [28], UADFV [29], Deepfake Detection Challenge Preview [30], and Deepfake Detection Challenge [31]. To ensure that the detector performance is not biased by specific pre-processing techniques, all datasets underwent uniform pre-processing. Additionally, any information that would identify the origin of the dataset has been eliminated, and the file names have been anonymized. For training, there are a total of 42,690 real images and 219,470 fake images. The validation dataset consists of 1,548 real images and 1,524 fake images. It is evident that there is a significant class imbalance present in the training data, as the ratio of fake and real images is approximately 5:1.
2.2 Data Preprocessing
2.2.1 Class Balancing for Model Training
One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset. Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, are trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
$\mathcal{D}_{i}=\mathcal{R}\cup\mathcal{F}_{i},\quad i\in\{1,2,3,4,5\}.$ (1)
We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [32]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, through the optimization:
$\theta_{i}=\arg\min_{\theta}\sum_{(x,y)\in\mathcal{D}_{i}}\mathcal{L}(M(x;\theta),y),$ (2)
where $\mathcal{L}$ represents the chosen loss function. After training on $\mathcal{D}_{i}$, the final weights $\theta_{i}$ from each stage are used to warm-start the next stage:
$\theta_{i+1}\leftarrow\theta_{i},\quad\text{for }i\in\{1,2,3,4\}.$ (3)
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images.
2.2.2 Data Augmentation and Normalization
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips(with 50% probability), random rotations(within $0^{o}-10^{o}$), color jittering (20% variation in brightness, contrast, saturation and hue), and random resized cropping where it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the DeiT architectures. For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain. Since all three models were pre-trained using the ImageNet dataset, the training images were normalized using the mean and standard deviation of the ImageNet dataset [32]. The input images were resized accordingly to the specific dimensions for the respective architectures.
2.3 Wavelet Feature Extraction for ConvNeXt
The wavelet transform is shown to be effective for image feature extraction as it decomposes images into multi-resolution frequency components [33]. The ability to decompose an image into different frequency sub-bands while preserving spatial relationships makes it particularly effective at identifying edges, textures, and subtle patterns of an image [34]. Using 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [23]. As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
Figure 1: Row- and column-wise filtering and downsampling to generate approximation and detail images through 2D-Discrete Wavelet Transform (DWT) using Haar wavelets.
The 2D-DWT is applied to an image $I(x,y)$ to decompose it into frequency sub-bands while preserving spatial information. Given an image $I\in\mathbb{R}^{M\times N}$, a single-level 2D-DWT using Haar wavelets generates four coefficient matrices: the approximate coefficients and three sets of detail coefficients (horizontal, vertical, and diagonal). The wavelet decomposition is performed using separable filtering along rows and columns. The Haar wavelet uses two filters:
  • 1.
    Low-pass filter ($\phi$): Coefficients $\left[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
  • 2.
    High-pass filter ($\psi$): Coefficients $\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
As shown in Figure 1, the sub-bands are computed as follows:
  1. 1.
    Row-wise filtering and downsampling:
    $L_{\text{row}}$ $=\left(I\ast\phi\right)\downarrow_{2,\text{cols}}$
    $H_{\text{row}}$ $=\left(I\ast\psi\right)\downarrow_{2,\text{cols}}$
  2. 2.
    Column-wise filtering and downsampling:
[…]
where $\ast$ indicates convolution; $\downarrow_{2,\text{cols}}$, and $\downarrow_{2,\text{rows}}$ indicate downsampling the columns and rows by 2, respectively; $A$ denotes low-frequency approximation coefficients; and $H,V,D$ denote high-frequency horizontal, vertical, and diagonal detail coefficients, respectively. For an RGB image $I_{r},I_{g},I_{b}$ consisting of three color channels, we apply the above transformation independently to each color channel:
$A_{c},H_{c},V_{c},D_{c}=\text{DWT}(I_{c}),\quad c\in\{r,g,b\},$ (4)
where $A_{c},H_{c},V_{c},D_{c}$ are the approximate and detail coefficients for each color channel $c$. The final feature representation is formed by concatenating the coefficients across all three channels:
$F=\text{concat}((A_{r},A_{g},A_{b}),(H_{r},H_{g},H_{b}),(V_{r},V_{g},V_{b}),(D_{r},D_{g},D_{b})).$ (5)
To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the ConvNeXt model. In Figure 2, we showed two such feature images with the approximate and detailed coefficients of two sample images from the dataset.
Figure 2: Real and Fake feature images after wavelet transform. The top left, top right, bottom left, and bottom right portions indicate the approximate, horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively. Different colormaps are applied to the horizontal, vertical, and diagonal detail coefficients, and their contrasts are increased for better visibility.
2.4 Details of the Model Architectures
Recall that our proposed solution utilizes three neural network models for a weighted ensemble approach for our final prediction. We chose each of these models to focus on different aspects of a deepfake image so that the ensemble approach provides better accuracy than individual models operating alone. The details of the models in our proposed solution are presented below.
2.4.1 EfficientNet
EfficientNet-B0 is a computationally efficient model whose architecture is optimized via neural architecture search (NAS) to maximize accuracy under a FLOP constraint, resulting in a highly efficient baseline that outperforms models with similar computational cost [21]. The core block of the EfficientNet architecture is the Mobile Inverted Bottleneck Convolution (MBConv) block inspired by MobileNetV2 [35]. The MBConv block consists of a 1x1 Convolution, which expands the input channels to increase feature dimensionality. Then, a 3x3 or 5x5 depthwise convolution is applied to capture spatial features efficiently. Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [36]. The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [37]. It is responsible for two main operations:
Squeeze: Global spatial information is aggregated using global average pooling to generate channel-wise descriptors. Given an input feature map $X\in\mathbb{R}^{H\times W\times C}$, the squeezed feature vector $z\in\mathbb{R}^{C}$ is obtained by squeezing function $F_{sq}(.)$ as:
$z_{c}=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}X_{c}(i,j),\quad c=1,2,...,C,$ (6)
where $X_{c}(i,j)$ represents the activation at spatial location $(i,j)$ in channel $c$.
Excitation: The descriptors generated from the Squeeze block are passed through a lightweight two-layer fully connected network with a bottleneck structure to model non-linear channel independencies.
After that, a skip connection adds the input to the output if the input and output shapes match. The swish activation [38] function is applied after each convolution except the final projection.
$\text{Swish}(x)=x\cdot\sigma(x)=\frac{x}{1+e^{-x}}$ (7)
EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage. A brief summary of different layers, the output shapes and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
Table 1: Summary of the EfficientNetB0 Model Architecture
Layer (type) Output Shape Param #
Conv2D + BatchNorm + SiLU [32, 112, 112] 928
DepthwiseSeparableConv [16, 112, 112] 1,448
InvertedResidual-1 $\times$ 2 [24, 56, 56] 16,714
InvertedResidual-2 $\times$ 2 [40, 28, 28] 46,640
InvertedResidual-3 $\times$ 3 [80,14,14] 242.930
InvertedResidual-4 $\times$ 3 [112, 14, 14] 543,148
InvertedResidual-5 $\times$ 4 [192, 7, 7] 2,026,348
InvertedResidual-6 [320, 7, 7] 717,232
Conv2d + BatchNorm + SiLU [1280, 7, 7] 412,160
AdaptiveAvgPool2d + Flatten [1280] 0
Linear [1] 1,281
Total Parameters 4,008,829
Trainable Parameters 4,008,829
Non-Trainable Parameters 0
2.4.2 Data Efficient Image Transformer (DeiT)
Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism. This allows transformers to learn long-range dependencies and global contexts of data [22]. The data-efficient image transformer, based on the backbone of the vision transformer, has been shown to work well in cases where training data is limited [39]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [40]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$, and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}$ (8)
In self-attention, the queries, keys, and values are derived from the same input sequence via linear transformations [41]. Multi-head self-attention (MSA) applies $h$ separate attention functions (heads) in parallel. Each head provides a sequence of outputs, which are concatenated and projected back to the original dimension. This allows the model to attend to different parts of the input sequence simultaneously.
DeiT is an effective alternative to traditional CNNs in learning global features [22], using knowledge distillation techniques and optimization strategies to improve generalization. DeiT learns from a pre-trained CNN model (that is, the RegNetY-16GF model [42]) and uses a distillation token during training, allowing it to learn faster and more efficiently with less data [39].
The DeiT architecture introduces a new hard distillation mechanism where they treat the hard decisions of the teacher model $y_{t}$ as true labels along with the actual true labels, $y$. If $Z_{t}$ are the logits of the teacher model and $Z_{s}$ are the logits of the student model, then the hard decision of the teacher model is defined as
$y_{t}=\text{argmax}_{c}(Z_{t}(c))$ (9)
Here, $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
$L^{\text{hard dist}}_{\text{global}}=\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y)+\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y_{t})$ (10)
The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$.
Our proposed solution utilizes the DeiT-B model, a 768-dimensional model with 12 heads and 86 million parameters, based on the ViT-B model [39]. The summary number of parameters and the output shapes of layers are shown in Table 2.
Table 2: Summary of the Data Efficient Image Transformer (DeiT) Model Architecture
Layer (type) Output Shape Param #
Conv2d [764, 14, 14] 590,592
Patch Embed [196, 768] 0
Dropout + Identity-1 [197, 768] 0
LayerNorm $\times$ 25 [197, 768] 38,400
Linear-1 $\times$ 12 [197, 2304] 21,261,312
Identity-2 [12, 197, 64] 0
(Linear-2 + Attention) $\times$ 12 [197, 768] 35,407,872
(Linear-3 + GELU) $\times$ 12 [197, 3072] 28,348,416
Dropout [197, 3072] 0
Mlp $\times$ 12 [197, 768] 0
Identity and Dropout [768] 0
Linear-4 [2] 1,538
Vision Transformer [2] 0
Total Parameters 85,648,130
Trainable Parameters 85,648,130
Non-Trainable Parameters 0
2.4.3 ConvNeXt Model
ConvNeXt combines the efficiency and inductive biases of traditional CNNs with the scalability and performance of transformers. It employs an inverted bottleneck block, similar to transformers’ MLP blocks, reducing FLOPs in downsampling layers. Here, batch normalization is replaced with layer normalization, aligning with transformer practices, and GELU is used instead of ReLU for smoother non-linearity [43].
Inspired by both ResNets [44] and Swin transformers [45], the architecture includes a patchify stem where input images are processed by a 4x4 convolution with stride 4, followed by layer normalization, to downsample the image into non-overlapping patches. This mimics ViT’s patch embedding but uses convolutions.
Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently. Then layer normalization is applied before and after the convolution to stabilize training. It normalizes features across the channel dimension for each spatial position. For feature $X[i,j,:]~{}\epsilon~{}\mathbf{R^{C}}$ the normalized output is:
$\hat{X}[i,j,c]=\frac{X[i,j,c]-\mu}{\sqrt{\sigma^{2}+\epsilon}}\cdot\gamma+\beta$ (11)
Here $\mu=\frac{1}{C}\sum_{c=1}^{C}X[i,j,c]$, $\sigma^{2}=\frac{1}{C}\sum_{c=1}^{C}(X[i,j,c]-\mu)^{2}$ , $\epsilon$ is a small constant, and $\gamma$ and $\beta$ are trainable parameters. This type of normalization works better than typical batch normalization [46]. After that, pointwise convolution is applied to expand (to 4x the input dimension) and contract the feature map, mimicking transformer MLP layers. GELU (Gaussian Error Linear Unit) [47] is applied instead of RELU for smoother non-linearity.
$\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}[1+\text{erf}(\frac{x}{\sqrt{2}})]$ (12)
A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [44]. Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with stride 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling. From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset. A summary of the different layers in the ConvNeXt-tiny architecture that we have implemented is presented in Table 3.
Table 3: Summary of the ConvNeXt Architecture
Layer Type Output Shape Param #
Conv2d + LayerNorm2d $[96,74,74]$ 4,896
CNBlock (×$\times 3$) $[96,74,74]$ 237,600
LayerNorm2d + Conv2d $[192,37,37]$ 74,112
CNBlock (×$\times 3$) $[192,37,37]$ 917,568
LayerNorm2d + Conv2d $[384,18,18]$ 295,680
CNBlock (×$\times 9$) $[384,18,18]$ 10,813,824
LayerNorm2d + Conv2d $[768,9,9]$ 1,181,184
CNBlock (×$\times 3$) $[768,9,9]$ 14,287,104
AdaptiveAvgPool2d + LayerNorm2d $[768,1,1]$ 1536
Flatten + Linear $[1]$ 769
Total Parameters 27,814,273
Trainable Parameters 27,814,273
Non-Trainable Parameters 0
2.4.4 Proposed Weighted Ensemble Architecture: CAE-Net
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search. The final architecture is shown in Figure 3.
Figure 3: The proposed CAE-Net architecture
2.5 Adversarial Perturbations
Adversarial perturbations are crafted noises or modifications added to classifier model inputs, often imperceptible to humans, that can significantly deteriorate model performance and lead to misclassification. Such adversarial perturbations can be universal or image-dependent [48] and are usually done in white-box settings where access to the model’s architecture and parameters is required. Also, adversarial attacks on a known network have been shown to work effectively against other networks as well [49], making black-box attacks an imminent threat for sensitive issues like deepfake detection [50]. We evaluate the robustness of our proposed models by testing them against a simple but effective adversarial perturbation, the Fast Sign Gradient Method (FGSM). FGSM is a white-box attack without specifying a target class. Using this method, adversarial images are created by calculating the gradient of loss for input data and adjusting the input by a small amount in the direction of the gradient sign [49].
$x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}J(\theta,x,y))$ (13)
Here $(x,y)$ are the original input image and its label, $\theta$ indicates model parameters, $x_{adv}$ is the adversarially perturbed image, and $\nabla_{x}J(\theta,x,y)$ is the gradient of the loss function $J$ with respect to the input image $x$. The parameter $\epsilon$ controls the magnitude of the perturbation added to the original image. This method works like a gradient ascent algorithm trying to maximize the loss with the smallest possible perturbation, $\epsilon$, which makes it barely noticeable to the human eye. Though FGSM is one of the simplest forms of perturbations, it has been shown to reduce the accuracy of high-quality detectors to a large extent [51]. Its effectiveness for binary classification problems makes it an ideal choice for testing the robustness of our proposed method.
3 Experimental Results
In this section, we demonstrate the performance of our proposed solution on the competition dataset. No prior work exists on this novel combined deepfake image detection dataset, which consists of images from eight different benchmark deepfake video datasets. To address this, we hypothesized that lightweight CNNs(e.g., EfficientNet) would excel in resource-constrained settings, modern CNNs (e.g., ConvNeXt) would leverage transformer-inspired designs for better feature extraction, and ViTs would capture the global context of images. We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset.
Table 4: Comparative Analysis of Different Model Variations
Classifier Models Accuracy
CNN models
  VGG-19 [52] 62.25%
  ResNet-50 [44] 69.00%
  MobileNet-V2 78.12%
  ResNet-34 [44] 80.32%
  RegNet [53] 80.92%
  Inception V3 [54] 84.38%
  ResNet-34 with CBAM 85.19%
  ResNet-34 with single SE 88.90%
  DenseNet-121 [55] 90.01%
  EfficientNet-B0 [21] 90.79%
  EfficientNet-B4 [21] 87.27%
  EfficientNet-B5 [21] 88.31%
  ResNeXt [56] 85.19%
  ConvNeXt [43] 88.18%
Frequency Domain Experimentations
  DCT with ResNet-34 71.48%
  ResNet-34 with Wavelet 80.44%
  EfficientNetB0 with Wavelet 82.48%
  Xception-Net [57] with Wavelet 87.76%
  ConvNeXt with Wavelet 89.49%
Study on Transformers
  ViT-B [58] 78.12%
  Swin Transformer [45] 86.52%
  MobileViT [59] 87.04%
  DeiT [39] 89.32%
Ensemble Results
  Maj. Voting (EfficientNet, DeiT, ConvNeXt) 94.24%
  Proposed Weighted Ensem. (EfficientNet, DeiT, ConvNeXt) 94.63%
Hyperparameters. For the training of our models, we have used the ADAM optimizer and binary categorical cross-entropy as our loss function. Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03385416s on a Tesla P100 GPU. The hyperparameters used for training are presented in Table 5.
Table 5: Hyperparameters for model training
Hyperparameters Details
Batch size 32
Number of Epochs in each session
   - EfficientNet 5
   - DeiT 5
   - ConvNeXt 4
Learning rate
   - Learning rate scheduler ReducedLROnPlateau
   - Initial learning rate 0.0001
   - Reduction factor 0.1
   - Patience 3
Result Analysis. We evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under the receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, real images were considered as class 0 and fake images as class 1. The performance of our model on these metrics is presented in Table 6 and the confusion matrix and the ROC curve are shown in Figure 4. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
Table 6: Performance of the Final Ensemble Architecture combining ConvNext, DeiT, and EfficientNet
Evaluation Metrics Performance
Accuracy 94.63%
Precision 0.937
Recall 0.956
F1 score 0.9464
Area under ROC 0.9737
Equal Error Rate 4.72%
Average Precision 0.9771
Figure 4: (a) Confusion matrix and (b) ROC curve for validation set
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection. In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. When the diverse insights from each model are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 6, showing the focus of irrelevant regions, leading to misclassifications.
(a) EfficientNet
(b) DeiT
(c) ConvNeXt
Figure 5: Importance map shown by Grad-CAM for different models
(a) EfficientNet
(b) DeiT
(c) ConvNeXt
Figure 6: Grad-CAM view of a few failed classifications of our proposed model
(a) EfficientNet (Untrained)
(b) DeiT (Untrained)
(c) ConvNeXt (Untrained)
(d) EfficientNet (Trained)
(e) DeiT (Trained)
(f) ConvNeXt (Trained)
Figure 7: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [60]. By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Fig. 7) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods. This highlights the generalization capability of our proposed ensemble model. After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Fig. 7(d), 7(e), and 7(f). There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets. A comparative analysis among existing works with our proposed architecture is shown in Table 7. Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [25]). We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is very poor for most architectures.
Table 7: Comparative Analysis of Average AUC Score of Different Models
[…]
DFD [27], DFDC [31], Celeb-DF [25]
Meso4 [63] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 75.9% 10.63%
DFD [27], DFDC [31], Celeb-DF [25]
FWA [64] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 82.1% 14.45%
DFD [27], DFDC [31], Celeb-DF [25]
Xception-C23 [26] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 86.4% 11.96%
DFD [27], DFDC [31], Celeb-DF [25]
DSP-FWA [25] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 87.4% 12.79%
DFD [27], DFDC [31], Celeb-DF [25]
Our Proposed Method Celeb-DF [25], DFDC [31], DFD [27], UADFV [29], 97.37% -
FaceShifter [28], DFDC Preview [30], FaceForensics++ [26]
Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradien Sign Method (FGSM). As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it. We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 8.
Table 8: Comparison of Our Proposed Architectures under Adversarial FGSM Attacks
Model Accuracy
$\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
EfficientNetB0 38.54% 46.48% 53.94%
Data Efficient Image Transformer 42.76% 21.94% 16.73%
ConvNeXt with wavelet 6.45% 6.51% 15.40%
Weighted Ensem.(EfficientNet, DeiT, ConvNeXt) 12.04% 13.48% 16.05%
Majority Voting (EfficientNet, DeiT, ConvNeXt) 25.52% 20.41% 19.76%
The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures. DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbation added to the image. Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks. If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly. However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying increased robustness against adversarial attacks for applications, such as deepfake detection.
4 Conclusion
In this work, we proposed an ensemble architecture for deepfake detection that captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training, and proposed a weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images. The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included Grad-CAM visualizations to map the focus regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy. An interesting future work could be incorporating lightweight architectures, reducing the computational overhead, and increasing the model’s adaptability to newer and sophisticated deepfake generation methods.